A mechatronic assembly is subjected to a final functional test. Suppose that defects occur at random in these assemblies, and that defects occur according to a Poisson distribution with parameter # 0.02. (a) What is the probability that an assembly will have exactly one defect? (b) What is the probability that an assembly will have one or more defects? (c) Suppose that you improve the process so that the occurrence rate of defects is cut in half to A 0.01. What effect does this have on the probability that an assembly will have one or 5. more defects?
Answer #1

Let X is a random variable shows the number of defects in the assembly. The pdf of X is

P(X = r)-e-0.02 0,02

(a)

The probability that assembly has exactly one defect is

0.020.02 i P(X = 1) = 0.01960 1!

(b)

The probability that assembly has exactly one or more defect is

e-0.020.020 01 0) = 0.01980

(c)

The probability that assembly has exactly one or more defect is

e-0010.01 01 = 0.00995